# ViEEG: Hierarchical Neural Coding with Cross-Modal Progressive Enhancement for EEG-Based Visual Decoding

### v2

@@ 5.3. Overall Performance
We conducted a series of ablation studies to assess the individual contributions of each core module in ViEEG, as summarized in Table 3. More detailed results are provided in the Appendix.
5.4.1. ablation on Hierarchical Embedding
This experiment analyzes the effectiveness of different levels of hierarchical EEG embeddings in supporting object recognition. Specifically, we evaluated configurations that utilized only one type of embedding at a time: $F_{b}$ (contour-level), $F_{f}$ (object-level), and $F_{r}$ (scene-level). In the subject-dependent setting, $F_{b}$ achieved 14.2% Top-1 and 41.8% Top-5 accuracy. In contrast, $F_{f}$ and $F_{r}$ both improved performance to approximately 22% Top-1 accuracy. These results suggest that while each representation captures distinct neural cues, none is sufficient alone to support optimal decoding performance. Notably, $F_{f}$ and $F_{r}$ embeddings, which encode mid- and high-level visual semantics, outperformed $F_{b}$, highlighting the greater salience of semantic and contextual cues in EEG decoding.
5.4.2. ablation on Cross Attention
We further examined the role of the cross-attention mechanism in integrating multi-level EEG features across the visual hierarchy. Ablating the cross-attention module (denoted as w/o C-Att) led to noticeable performance degradation across all settings. For instance, in the subject-dependent setting, the Top-1 and Top-5 accuracies dropped from 34.1% and 71.3% to 32.4% and 66.8%, respectively, when cross-attention was removed. In the subject-independent setting, similar trends were observed, with the full ViEEG model outperforming the variant without cross-attention by notable margins. These findings underscore the importance of cross-attention for aligning and integrating hierarchical EEG embeddings, thereby enhancing overall model robustness and generalization across subjects. The performance gain is particularly evident in the integration of $F_{f}$ and $F_{r}$, as visualized in Figure 4, where the majority of subjects demonstrate that adding cross-attention for hierarchical visual information integration effectively improves EEG decoding of the integrated features.
@@
Figure 4. Impact of Cross-Attention on Feature Integration. Subfigures 1 and 2 show the Top-1 and Top-5 accuracies of foreground objcet embedding ($F_{f}$) with and without cross-attention, and subfigures 3 and 4 report the same metrics for scene embedding ($F_{r}$).
6. Discussion
6.1. Representational Analysis
To assess the representational consistency of ViEEG, we visualize the Representational Similarity Matrix (RSM) in the form of heatmaps, using data from Subject 8 of the THINGS-EEG dataset. The RSM quantifies pairwise similarity among neural representations across different visual categories, offering insight into how well the learned embeddings preserve intrinsic EEG structure. For comparison, we also visualize the corresponding RSMs generated by NICE and ATM, as shown in Figure 5. More results from other subjects are provided in the Appendix.
The RSMs of all three methods exhibit block-diagonal patterns when test samples are ordered by category, indicating basic inter-class separability. However, ViEEG distinguishes itself by preserving sharper main diagonals within each block, reflecting stronger intra-class consistency and more precise instance-level matching. For example, in the ”Food” category, ViEEG consistently localizes target representations with clear diagonal peaks, while NICE and ATM produce more diffuse patterns with less focus on the diagonal. This suggests that ViEEG can better discriminate among similar samples within a class. In contrast, ambiguous categories like ”Tool” and ”Sports” naturally show less defined separability, and the “Other” class presents no clear structure due to its heterogeneous nature. This dual preservation of class-level structure and instance-level discrimination highlights ViEEG’s potential in real-world EEG retrieval and classification tasks, where both categorical understanding and exemplar precision are essential.
Figure 5. Representational Similarity Matrices of ViEEG, NICE, and ATM. Top row: RSMs across six major categories (Animal, Food, Vehicle, Tool, Sports, and Other). Bottom row: Zoomed-in view of the Food category.
6.2. Brain Decoding Results Analysis
6.2.1. Zero-shot Image Retrieval
@@
Figure 6. Zero-shot Image Retrieval with Different Embeddings (Subject 8). Left: Basketball example shows BOM excels due to clean contours. Right: Aircraft carrier case favors FO due to complex shape and background blending.
To further evaluate ViEEG’s capacity for visual concept decoding from EEG, we conduct zero-shot image retrieval experiments based on different visual embeddings: binary object mask (BOM), foreground object (FO), and raw scene (RS) embedding, as well as concatenated triple visual embedding. As shown in Figure 6, each embedding variant demonstrates distinct advantages depending on the object’s visual structure and contextual complexity. For example, for the basketball, BOM yields the highest retrieval accuracy, likely due to the clean, circular silhouette of the ball. In contrast, FO and RS embeddings underperform, possibly due to background clutter or inconsistent attentional focus during EEG acquisition. Conversely, for the aircraft carrier, the complex object geometry and blended ocean background reduce BOM and RS effectiveness, making FO the most discriminative embedding.
Crucially, concatenating all three embeddings—forming the Tri-View representation—consistently yields superior performance. This suggests a complementary mechanism, where confident predictions from any individual view elevate the joint similarity score, mitigating retrieval errors caused by weaker modalities. Additional retrieval results are provided in the Appendix to further demonstrate the robustness and reliability of our framework.
Figure 7. Parameter Analysis for the Attention Module. Top row: Accuracy trends for varying numbers of attention layers; Bottom row: Accuracy trends for different numbers of attention heads. Both Top-1 and Top-5 metrics under subject-dependent and subject-independent settings are reported.
6.2.2. Image Reconstruction
@@
To further assess the visual fidelity of ViEEG’s learned representations, we perform image reconstruction using visual embeddings derived from EEG. As shown in Figure 1, the reconstructed images preserve key semantic information of the original stimuli, including object category, shape, and spatial structure. Notably, ViEEG is capable of reconstructing both the foreground object and its contour, demonstrating its capacity to capture fine-grained visual features from brain signals. Due to space limitations, more reconstruction examples and implementation details are provided in the Appendix.
6.3. Parameter Analysis
ViEEG integrates a cross-attention module to hierarchically fuse contour, object, and context-level EEG features. We perform grid search experiments to explore the impact of attention layers and heads under both subject-dependent and subject-independent settings (Figure 7). Varying the number of attention layers from 0 to 5, we find that a single-layer configuration yields the best Top-1 accuracy in both settings. Deeper attention stacks lead to overfitting or marginal gains, making one layer the optimal balance between performance and complexity. For attention heads, we test 1 to 6 heads. Three heads achieve the highest accuracy in the subject-dependent case, while subject-independent performance remains relatively stable across configurations. Overall, ViEEG performs best with one attention layer and three heads.
7. Conclusion
@@
In this work, we proposed ViEEG, a novel EEG-based visual decoding framework that emulates the hierarchical structure of human visual perception. By integrating biologically motivated image decomposition with hierarchical EEG encoding and cross-attention fusion, ViEEG captures the progressive flow of visual information from edge detection to semantic and contextual understanding. This design directly addresses the critical issue of hierarchical neural encoded neglect (HNEN) in conventional EEG decoding approaches. Through extensive experiments on the THINGS-EEG dataset, ViEEG significantly outperformed state-of-the-art baselines in both subject-dependent and cross-subject zero-shot recognition tasks. Ablation studies further validate the importance of each component, particularly the hierarchical embedding structure and cross-attention integration. Beyond performance gains, ViEEG offers a cognitively plausible model that strengthens the connection between neuroscience and artificial intelligence. We believe this work opens up new possibilities for neuro-inspired learning and brings us closer to more accurate, interpretable, and generalizable brain-computer interface systems.
@@
Appendix A Details of the Datasets
@@
The THINGS-EEG dataset (Gifford et al., 2022), used in this study, is a large-scale EEG dataset designed to model visual object recognition. It includes data from ten participants who were exposed to images from the THINGS database. The images were presented using a rapid serial visual presentation (RSVP) paradigm, where each image was shown for 100 ms followed by a blank screen for 100 ms, maintaining a stimulus onset asynchrony (SOA) of 200 ms. The dataset contains 16,540 training image conditions (1,654 concepts, each with 10 images, repeated four times) and 200 test image conditions (200 concepts, each with one image, repeated once). The THINGS image dataset comprises six major categories, including animals (e.g., cats, dogs), vehicles (e.g., airplanes, ships), food items (e.g., cake, corn), tools (e.g., cameras, phones), sports equipment (e.g., balls, golf clubs), and other miscellaneous categories.
Figure 8. The corresponding EEG 63-electrode placement of THINGS-EEG datasets.
EEG signals were recorded using a 64-channel cap according to the 10-10 system, sampled at 1000 Hz, and filtered within a frequency range of 0.1–100 Hz. Preprocessing included segmenting the data into epochs from 200 ms before to 800 ms after stimulus onset, baseline correction using the mean pre-stimulus interval, and down-sampling to 100 Hz. To reduce noise, repeated trials were averaged, and multivariate noise normalization was applied using training data. For input into the image encoder, images were resized to $224\times 224$ pixels and normalized.
Appendix B Comparison Methods
@@
To validate the efficacy of our proposed approach, we compare ViEEG with a range of classical and recent methods spanning EEG cognitive decoding and visual decoding. Below, we provide detailed descriptions of the competing approaches:
@@
  • BraVL (Du et al., 2023): BraVL employs multimodal learning to jointly model brain, visual, and linguistic features, maximizing intra- and inter-modality mutual information to improve consistency and data efficiency.
@@
  • LSTM (Hochreiter, 1997): Long Short-Term Memory networks are leveraged to capture long-term temporal dependencies in sequential EEG data.
@@
  • ConvNet (Schirrmeister et al., 2017): A convolutional neural network designed for end-to-end EEG decoding, focusing on extracting spatial and spectral features directly from raw EEG signals.
@@
  • Conformer (Song et al., 2022): This architecture combines convolutional neural networks with Transformer modules to extract both local and global dependencies in sequential EEG data.
@@
  • DGCNN (Song et al., 2018): The Dynamical Graph CNN employs a learnable adjacency matrix along with Chebyshev filters to capture dynamic relationships in EEG data for emotion classification.
@@
  • EEGNet (Lawhern et al., 2018): A lightweight convolutional network that efficiently extracts temporal and spatial features from EEG signals, optimized for BCI applications.
@@
  • MB2C (Wei et al., 2024): The Multimodal Bidirectional Cycle Consistency framework utilizes dual-GAN architectures to generate and reconcile modality-specific features, effectively bridging the modality gap in zero-shot tasks, EEG classification, and image reconstruction.
@@
  • NICE (Song et al., 2024): A self-supervised framework that learns image representations from EEG signals by incorporating attention modules to capture spatial correlations within brain activity.
@@
  • ATM (Li et al., 2024): The Adaptive Thinking Mapper projects neural signals into a shared subspace with CLIP embeddings via a two-stage EEG-to-image generation strategy, enabling zero-shot visual decoding and reconstruction.
@@
  • CognitionCapturer (CogCap) (Zhang et al., 2024): This unified framework enhances EEG signal representations by leveraging multimodal data and modality-specific expert encoders. It employs a diffusion prior to map EEG embeddings to the CLIP space, achieving high-fidelity reconstruction of visual stimuli without requiring fine-tuning.
@@
Appendix C Whole Results of ablation Study
@@
Table 5. Object Recognition Accuracy for Different Hierarchical Feature Combinations in ViEEG with cross-attention module
@@
Table 6. Object Recognition Accuracy for Different Hierarchical Feature Combinations in ViEEG without cross-attention module
Due to space limitations, only partial information is presented in the main text. The complete results of the ablation experiments on Hierarchical Embedding and cross-attention are provided here. Figure 5 shows the object recognition accuracy of Hierarchical Embedding in the full ViEEG model, while Figure 6 presents the object recognition accuracy of Hierarchical Embedding in the ViEEG model without cross-attention.
C.1. ablation on Hierarchical Embedding
The results presented in Table 5 showcase the performance of different hierarchical EEG feature combinations on object recognition. These results were obtained from the ViEEG model with the full cross-attention mechanism.

### v3

@@ 4.4 Ablation Study
We conducted ablation study to assess the contributions of each core module in ViEEG, as summarized in Table 3.
@@
Figure 4: Impact of cross-attention integration. Subfigures 1&2 show Top-1/5 accuracy of object embedding ($F_{f}$) with/without attention, and subfigures 3&4 show the same metrics for context embedding ($F_{r}$).
Ablation on Hierarchical Embedding.
This experiment analyzes the effectiveness of different levels of hierarchical EEG embeddings. Specifically, we evaluated configurations that utilized only one type of embedding at a time: $F_{b}$ (contour-level), $F_{f}$ (object-level), and $F_{r}$ (scene-level). In the subject-dependent setting, $F_{b}$ achieved 14.2% Top-1 and 41.8% Top-5 accuracy. In contrast, $F_{f}$ and $F_{r}$ both improved performance to approximately 22% Top-1 accuracy. These results suggest that while each representation captures distinct neural elements, none is sufficient alone to support optimal decoding performance. Notably, $F_{f}$ and $F_{r}$ embeddings, which encode mid-level and high-level visual semantics, outperformed $F_{b}$, highlighting the greater salience of semantic and contextual elements in EEG brain decoding.
Ablation on Cross Attention.
We further examined the role of cross-attention mechanism in integrating multi-level EEG features across the visual hierarchy. Ablating the cross-attention led to noticeable performance degradation. For instance, in the subject-dependent setting, the Top-1 and Top-5 accuracies dropped from 34.1% and 71.3% to 32.4% and 66.8%, respectively, when cross-attention was removed. In the subject-independent setting, similar trends were observed, with the full ViEEG outperforming the variant without cross-attention by notable margins. These findings underscore the importance of cross-attention for aligning and integrating hierarchical EEG embeddings, thereby enhancing overall model robustness and generalization across subjects. The performance gain is particularly evident in the integration of $F_{f}$ and $F_{r}$, as visualized in Figure 4, where the majority of subjects demonstrate that adding cross-attention for hierarchical visual information integration effectively improves EEG decoding of the integrated features.
@@
Figure 5: Zero-shot image retrieval with different embeddings (Subject 8). Left: basketball example shows BOM excels due to clean contours. Right: aircraft carrier case favors FO due to complex shape and background blending.
4.5 Representational Analysis
Representational Similarity Matrices.
To assess the representational consistency of ViEEG, we visualize the representational similarity matrices (RSMs) as heatmaps using Subject 8 from the THINGS-EEG dataset. RSMs reflect pairwise similarities among neural embeddings across visual categories, providing insights into how well each model captures the intrinsic EEG structure. For comparison, RSMs from NICE and ATM are also shown in Figure 6, with additional subjects in supplementary Appendix.
All three models exhibit block-diagonal patterns when test samples are ordered by category, indicating basic class-level separation. However, ViEEG shows noticeably sharper diagonals within each block, reflecting stronger intra-class consistency and more accurate instance-level alignment. For example, in the “Food” category, ViEEG maintains focused diagonal peaks, whereas NICE and ATM produce more diffuse patterns. This suggests ViEEG achieves finer discrimination within categories. In less structured or ambiguous classes such as “Tool,” “Sports,” and “Other,” all models show reduced separability, though ViEEG retains relatively clearer patterns. Overall, ViEEG better preserves both category structure and exemplar specificity, which is crucial for downstream EEG retrieval and classification tasks.
Figure 6: Representational similarity matrices (RSM) across categories (Animal, Food, Vehicle, Tool, Sports, and Others), and zoomed-in view of Food category.
Zero-shot Image Retrieval.
We further evaluate ViEEG’s decoding ability through zero-shot image retrieval using different visual embeddings: binary object mask (BOM), foreground object (FO), raw scene (RS), and their concatenation (Triple). As shown in Figure 5, each embedding performs differently based on the visual characteristics of the object. For instance, BOM performs best on simple silhouettes like basketballs, while FO excels for complex scenes like aircraft carriers where background blending challenges other embeddings. Notably, the triple fusion consistently achieves the highest accuracy, leveraging complementary information from all views. This fusion mitigates the weaknesses of individual embeddings and enhances retrieval robustness. Additional examples are provided in supplementary Appendix.
Image Reconstruction.
@@ 4.5 Representational Analysis
To further assess the visual fidelity of ViEEG’s learned representations, we perform image reconstruction using visual embeddings derived from EEG. As shown in Figure 2, the reconstructed images preserve key semantic information of the original stimuli, including object category, shape, and spatial structure. Notably, ViEEG is capable of reconstructing the foreground object and contour, demonstrating its capacity to capture fine-grained visual features from brain signals. Due to space limitations, more reconstruction examples and implementation details are provided in supplementary Appendix.
5 Conclusion
@@
In this work, we proposed ViEEG, a novel EEG visual decoding framework that emulates the hierarchical structure of human visual perception. By integrating biologically motivated image decomposition with hierarchical EEG encoding and cross-attention fusion, ViEEG captures the progressive flow of visual information from edge detection to semantic and contextual understanding. This design directly addresses the critical issue of hierarchical neural encoded neglect in conventional EEG decoding approaches. Through extensive experiments on the THINGS-EEG and THINGS-MEG datasets, ViEEG significantly outperformed state-of-the-art baselines in both subject-dependent and subject-independent zero-shot recognition. Ablation study further validate the importance of each component. Beyond performance gains, ViEEG offers a cognitively plausible model that strengthens the connection between neuroscience and artificial intelligence. We believe our work advances neuro-inspired learning toward more robust, interpretable, and generalizable BCI systems.
@@
6 Details of the Datasets
6.1 THINGS-EEG
We conducted experiments on the large-scale THINGS-EEG dataset (Gifford et al. 2022), designed to support neural decoding of visual object recognition. The dataset contains EEG recordings from ten healthy adult participants (mean age: 28.5 years; 8 female, 2 male), all with normal or corrected-to-normal vision. EEG signals were acquired using a 64-channel (see Figure 7) cap arranged according to the 10–10 international system and sampled at 1000 Hz, with online band-pass filtering between 0.1–100 Hz and referencing to Fz.
Figure 7: The corresponding EEG electrode placement of THINGS-EEG dataset.
@@ 6.1 THINGS-EEG
Visual stimuli were selected from the THINGS image database (Hebart et al. 2019), which includes naturalistic object images spanning 1854 distinct concepts grouped into 27 higher-level categories. The THINGS image dataset comprises six major categories, including animals (e.g., cats, dogs), vehicles (e.g., airplanes, ships), food items (e.g., cake, corn), tools (e.g., cameras, phones), sports equipment (e.g., balls, golf clubs), and other miscellaneous categories. For model training and evaluation, these concepts were split into 1654 training and 200 testing categories. Each training concept was associated with ten images, while each test concept had one image. The experiment followed a rapid serial visual presentation (RSVP) paradigm (Thorpe, Fize, and Marlot 1996), where participants viewed 20 images per trial, each presented for 100 ms followed by a 100 ms blank screen (i.e., 200 ms SOA). An orthogonal target detection task was used to maintain attention.
Each participant completed four sessions, resulting in a total of 82,160 trials per subject, including 16,540 training image trials (each repeated four times) and 200 test images (each repeated 80 times). Trials containing target stimuli were excluded. EEG data were segmented from 200 ms before to 800 ms after stimulus onset, followed by baseline correction and downsampling to 250 Hz. All 64 channels were retained, and signals were averaged across repeated presentations of the same image to enhance signal-to-noise ratio. Following prior work (Song et al. 2024), multivariate noise normalization was applied to the training set. Input images were resized to 224×224 pixels and normalized before being fed into the visual encoder.
6.2 THINGS-MEG
We additionally evaluated our model on the THINGS-MEG dataset (Hebart et al. 2023), which provides magnetoencephalography (MEG) recordings from four participants across 12 sessions. Visual stimuli were drawn from the THINGS database (Hebart et al. 2019), encompassing 1854 object concepts and over 26,000 curated naturalistic images. During each session, participants viewed images for 500 ms followed by a blank interval of 1000 ± 200 ms while maintaining central fixation. An orthogonal oddball detection task was employed to ensure sustained attention.
For zero-shot evaluation, 200 object concepts were held out from the training set. The training set consisted of 1854 concepts, each paired with 12 unique images (1854 × 12 × 1), and the test set included 200 concepts with one image repeated 12 times (200 × 1 × 12). MEG data were acquired with a 271-channel whole-head system and epoched from 0 to 1000 ms after stimulus onset. Preprocessing involved band-pass filtering in the 0.1–100 Hz range, baseline correction, and downsampling to 200 Hz. To improve signal stability, we averaged repeated responses for each image. Due to the limited number of participants, statistical analysis was not conducted on this dataset, and it serves as a supplementary validation of the EEG-based findings.
Subject 1 Subject 2 Subject 3 Subject 4 Ave
Method top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
Subject dependent - train and test on one subject
Conformer (Song et al. 2022) 9.1 27.1 20.8 47.4 15.2 40.7 9.8 27.9 13.7 35.8
DGCNN (Song et al. 2018) 6.7 22.0 14.1 36.2 11.9 35.8 7.7 22.8 10.1 29.2
EEGNet (Lawhern et al. 2018) 9.8 29.2 17.9 48.9 14.8 41.3 9.1 28.7 12.9 37.0
Mb2C (Wei et al. 2024) 9.3 33.6 20.6 49.2 18.2 44.3 10.2 33.6 14.6 39.9
NICE (Song et al. 2024) 11.5 35.6 25.7 54.4 21.0 47.8 11.2 35.2 17.4 43.3
ATM (Li et al. 2024) 8.0 29.3 30.2 61.5 20.3 50.5 11.8 33.3 17.6 43.7
ViEEG 16.6 50.3 37.4 79.1 30.6 73.9 17.2 49.5 25.5 63.2
Table 4: Overall Top-1/5 accuracy (%) on THINGS-MEG for 200-way zero-shot recognition.
7 Comparison Methods
@@
To validate the efficacy of our proposed approach, we compare ViEEG with some classical and recent methods for EEG cognitive decoding and visual decoding. Below, we provide detailed descriptions of the baseline approaches:
@@
  • BraVL [IEEE TPAMI’2023] (Du et al. 2023): BraVL employs multimodal learning to jointly model brain, visual, and linguistic features, maximizing intra- and inter-modality mutual information to improve consistency and data efficiency.
@@
  • LSTM (Hochreiter 1997): Long Short-Term Memory networks are leveraged to capture long-term temporal dependencies in sequential EEG data.
@@
  • ConvNet (Schirrmeister et al. 2017): A convolutional neural network designed for end-to-end EEG decoding, focusing on extracting spatial and spectral features directly from raw EEG signals.
@@
  • Conformer [IEEE TNSRE’2022] (Song et al. 2022): This architecture combines convolutional neural networks with Transformer modules to extract both local and global dependencies in sequential EEG data.
@@
  • DGCNN [IEEE TAFFC’2018] (Song et al. 2018): The Dynamical Graph CNN employs a learnable adjacency matrix along with Chebyshev filters to capture dynamic relationships for EEG emotion classification.
@@
  • EEGNet (Lawhern et al. 2018): A lightweight convolutional network that efficiently extracts temporal and spatial features from EEG signals, optimized for BCI applications.
@@
  • MB2C [ACM MM’2024] (Wei et al. 2024): The Multimodal Bidirectional Cycle Consistency framework utilizes dual-GAN architectures to generate and reconcile modality-specific features, effectively bridging the modality gap in zero-shot tasks, EEG classification, and image reconstruction.
@@
  • NICE [ICLR’2024] (Song et al. 2024): A self-supervised framework that learns image representations from EEG by incorporating attention modules to capture spatial correlations within brain activity.
@@
  • ATM [NeurIPS’2024] (Li et al. 2024): The Adaptive Thinking Mapper projects neural signals into a shared subspace with CLIP embeddings via a two-stage EEG-to-image generation strategy, enabling zero-shot visual decoding and reconstruction.
@@
  • CognitionCapturer (CogCap) [AAAI’2025] (Zhang et al. 2024): This unified framework enhances EEG signal representations by leveraging multimodal data and modality-specific expert encoders. It employs a diffusion prior to map EEG embeddings to the CLIP space, achieving high-fidelity reconstruction of visual stimuli without requiring fine-tuning.
@@
8 Experiment in THINGS-MEG Dataset
Due to the limited number of subjects in the THINGS-MEG dataset, we conducted subject-dependent experiments, training and testing separately on each subject. Table 4 reports the 200-way zero-shot recognition results. As shown in Table 4, ViEEG also achieves superior performance on the THINGS-MEG dataset, outperforming all baselines in both top-1 and top-5 accuracy across subjects. These results demonstrate that our neuroscience-inspired framework is not only effective for EEG-based decoding, but also generalizes well to other neural signals such as MEG, further highlighting the robustness and cross-modality potential of our proposed ViEEG.
9 Whole Results of Ablation Study and Parameter Analysis
@@
Table 5: Object recognition accuracy for different hierarchical representation integration in ViEEG with cross-attention module
@@
Table 6: Object recognition accuracy for different hierarchical representation integration in ViEEG without cross-attention module
Figure 8: Detailed ablation study of attention on hierarchical visual embeddings in ViEEG.
Due to space limitations, only partial information is presented in the main text. The complete results of the ablation experiments on Hierarchical Embedding and cross-attention are provided here. Figure 5 shows the object recognition accuracy of Hierarchical Embedding in the full ViEEG, while Figure 6 presents the object recognition accuracy of Hierarchical Embedding in ViEEG without cross-attention.
9.1 Ablation on Hierarchical Embedding
The results presented in Table 5 showcase the performance of different hierarchical EEG feature combinations on object recognition. These results were obtained from ViEEG with the full cross-attention mechanism.